OpenStudy (anonymous):

Which polynomial has a graph that passes through the given points?(-3,10), (-1,0)

OpenStudy (anonymous):

I would try to find the slope first. So the formula for slope is (y2-y1)/(x2-x1) For your numbers i get (0-10)/(-1--3)=-5 Then i use the form y=mx+b (you can do this another way, just my suggestion) m is your slop of -5 You can plug in either of the two points to solve for b, it is usually easier to use the one with a zero 0=-5(-1)+b b=-6 Formula y=-5x-6